OpenStudy (anonymous):

The radius of convergence for x^2 * e^x ?

OpenStudy (freckles):

was there suppose to be a series thing here ?

OpenStudy (anonymous):

It said "Find the first five non-zero terms of Taylor series centered at x=0 for the function below."

OpenStudy (freckles):

and for that part of the question can you tell me what you got ?

OpenStudy (anonymous):

x^2 + x^3 + x^4 / 2 + x^5 / 6 + x^6/24

OpenStudy (freckles):

\[\sum_{n=1}^{\infty}\frac{x^{n+1}}{(n-1)!}\]

OpenStudy (freckles):

I would suggest using the ratio test

OpenStudy (anonymous):

Okie

OpenStudy (anonymous):

The radius of convergence is infinity.

OpenStudy (freckles):

why do you say that?

OpenStudy (freckles):

can you tell me what have simplified this part to: \[a_{n+1} \cdot \frac{1}{a_n}=\frac{x^{n+2}}{n!} \cdot \frac{(n-1)!}{x^{n+1}}\]

OpenStudy (freckles):

anyways good work